# Inference for Heterogeneous Treatment Effects with Efficient Instruments and Machine Learning

### v1

@@ 4.3.2 Effect of Education on Wage
MLIV (random forest) 0.14 0.08 $[-0.01,0.3]$ $[0.01,0.60]$
@@
Also for the Card dataset, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV) to see if we can detect heterogeneity of the treatment effect. As the variable $A_{i}$, we take the years of experience of individual $i$. The estimated heterogeneous treatment effects with respect to experience are displayed in Figure 5. Again, we only show the results for the machine learning methods gam and xgboost and omit the results for random forest since they look similar. Moreover, we still use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing.
@@ 5 Discussion
A limitation of our approach is that it only considers heterogeneity with respect to a univariate quantity, i.e. the dimension of $A_{i}$ is $1$. If the treatment effect is potentially heterogeneous with respect to the full set of covariates $X_{i}$, i.e. the true model is $Y_{i}=\beta(X_{i})D_{i}+g(X_{i})+\epsilon_{i}$, equation (22) identifies $\mathbb{E}[\beta(X_{i})w(Z_{i},X_{i},a)|A_{i}=a]$ with the weight function
$w(Z_{i},X_{i},a)=\frac{(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}}{\mathbb{E}[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}|A_{i}=a]}.$
In particular, this is not just the projection $\mathbb{E}[\beta(X_{i})|A_{i}=a]$ of the true heterogeneous treatment effect on $A_{i}$.
A straightforward extension of our approach would be to use multivariate kernel smoothing to deal with multivariate $A_{i}$ and the results should carry over. However, note that the conditions on the convergence rates of the nuisance function estimators (Assumption 5) depend on the bandwidth $h$. In higher dimensions, they will depend on $h^{d}$, where $h$ is the dimension of $h$ and it will quickly become unrealistic to have machine learning methods that achieve these convergence rates.
We can also take one step back. In principle, we can replace the outer conditional expectations in (22) by arbitrary machine learning methods potentially also allowing for higher-dimensional $A_{i}$. A natural assumption might be that $A_{i}=X_{i}$ which is potentially high-dimensional, but that the treatment effect is only heterogeneous with respect to a small subset of the components of $X_{i}$, so one could use methods enforcing sparsity for the outer conditional expectation. If the machine learning method in the numerator allows for inference and the machine learning method in the denominator is consistent, one might still be able to perform inference for the heterogeneous treatment effect.
Moreover, we have only provided pointwise confidence intervals for our method. However, it would also be interesting to have uniform confidence bands. We expect that this is possible using multiplier bootstrap, as for example done in [17].
@@ Acknowledgements
CS received funding from the Swiss National Science Foundation, grant no. 214865. ZG acknowledges financial support for visiting the Institute of Mathematical Research (FIM) at ETH Zurich. ZG was partly supported by NIH R01LM013614, NIH R01AG086379, and NSF DMS 2413107.
@@ Appendix A Technical Conditions for the Heterogeneous Treatment Effect
The following assumptions are needed to be able to estimate and perform inference for the heterogeneous treatment effect $\beta(a)$.
@@
    The density $p_{A}(\cdot)$ of $A_{i}$ with respect to Lebesgue measure satisfies that $p_{A}(a)\geq C$ for a constant $C>0$ independent of $N$.
@@
    $\sigma_{Q}^{2}(a)\asymp\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a]$.
@@
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=a\right]\lesssim\sigma_{Q}^{2}(a)$.
@@
    The function $\rho(x)=(\beta(x)-\beta(a))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}(a)<C$.
@@
    The function $\rho(x)=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(a)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(a))\mathbb{E}[\epsilon_{i}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(a)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(a))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(a)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(a))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
@@
    The function $\rho(x)=\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(a)<C$.
@@
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(a)<C$.
@@
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|A_{i}=x]p_{A}(x)$ is twice differentiable for $x$ in the support of $A_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
@@
    $\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{A_{i}-a}{h}\right)\right|^{2+\eta}\right]}{\sigma_{Q}(A)^{2+\eta}}\leq Ch.$
@@
    $\frac{\mathbb{E}\left[\left(\epsilon_{i}+(\beta(A_{i})-\beta(a))(D_{i}-\phi(X_{i}))\right)^{4}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{4}K\left(\frac{A_{i}-a}{h}\right)^{4}\right]}{\sigma_{Q}^{4}(a)}\ll Nh^{2}.$
@@
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{4}K\left(\frac{A_{i}-a}{h}\right)^{2}\right]=O(h).$
@@
Note that all the quantities in Assumption 6 may depend on $N$ even if it is not reflected in the notation. Assertion 1 intuitively means that we need to have enough samples $i$ with $A_{i}$ close to $a$ to be able to estimate the heterogeneous treatment effect. Assertions 2 and 3 are needed in the case where the IV strength $\sigma_{Q}^{2}(a)$ is decreasing. Essentially they state that a decreasing $\sigma_{Q}^{2}(a)$ is attributed to $f(Z_{i},X_{i})$ being close to $\phi(X_{i})$. They are for example satisfied if $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ and $\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|Z_{i},X_{i}]$ are both bounded away from zero and infinity independent of $N$. Assertions 4-11 are smoothness assumptions on $\beta(x)$, $p_{A}(x)$ and conditional moments of $\epsilon_{i}$, $D_{i}$, $f(Z_{i},X_{i})$ and $\phi(X_{i})$. These smoothness conditions need to hold independently of $N$. In view of Lemma 24 in Appendix D.5 (with the kernel $K$ potentially replaced by $K^{2+\eta}$ or $K^{4}$), assertions 12-14 are comparably weak and hold in particular, if the quantities inside the expectations satisfy similar smoothness conditions as assertions 4-11.
@@ C.1 Explicit Expression for the Robust Confidence Set
Using basic manipulations, we can derive explicit expressions for $\mathcal{C}_{N}^{\alpha}(a)$. For this, define the quantities
$A_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)$
$B_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{A_{i}-a}{h}\right)$
$C_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$E_{N}$ $=\frac{2}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{D,i}^{k}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}$
$F_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{A_{i}-a}{h}\right)^{2}.$
@@
$\mathcal{C}_{N}^{\alpha}(a)=\left\{\gamma:(A_{N}-\gamma B_{N})^{2}\leq\frac{z_{1-\alpha/2}^{2}}{Nh}\left(C_{N}-\gamma E_{N}+\gamma^{2}F_{N}-h(A_{N}-\gamma B_{N})^{2}\right)\right\}$
or equivalently $\mathcal{C}_{N}^{\alpha}(a)=\left\{\gamma:R_{N}\gamma^{2}+S_{N}\gamma+T_{N}\leq 0\right\}$ with
@@
$\mathcal{C}_{N}^{\alpha}(a)=\begin{cases}\left[\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}>0\\ \mathbb{R}\setminus\left[\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}<0.\end{cases}$ (27)
By construction, we have that $\hat{\beta}(a)=A_{N}/B_{N}\in\mathcal{C}_{N}^{\alpha}(a)$. Hence, in the case $R_{N}>0$, it will always hold that $S_{N}^{2}-4R_{N}T_{N}>0$. In the case $R_{N}<0$, it may happen that $S_{N}^{2}-4R_{N}T_{N}<0$, in which case, $\mathcal{C}_{N}^{\alpha}(a)=\mathbb{R}$.

### v2

@@ 4.3.2 Effect of Education on Wage
MLIV (random forest) 0.14 0.08 [0.01,0.3][-0.01,0.3] $[0.01,0.60]$
@@
Also for the Card dataset, we apply the methods (het. T.E. mlIV) and (het. T.E. linearIV) to see if we can detect heterogeneity of the treatment effect. As the variable $V_{i}$, we take the years of experience of individual $i$. The estimated heterogeneous treatment effects with respect to experience are displayed in Figure 5. Again, we only show the results for the machine learning methods gam and xgboost and omit the results for random forest since they look similar. Moreover, we still use the Epanechnikov kernel and the bandwidth according to the normal reference rule as well as a smaller bandwidth to ensure undersmoothing.
@@ 5 Discussion
An important avenue for further research are methods to decide from data if a constant treatment effect model is enough or if a heterogeneous treatment effect model is more appropriate. As an informal test, one can check if the homogeneous treatment effect estimate (i.e., a horizontal line) lies inside the confidence sets for the heterogeneous treatment effect. In fact, from the corresponding plots in our two real-data case studies, we observed that this is (more or less) the case which leads to the conclusion that the heterogeneous treatment effect model is probably not needed. However, the confidence sets that we provided are only pointwise. Hence, it would also be interesting to have uniform confidence bands. We expect that this is possible using multiplier bootstrap, as for example done in [24].
[…]
A limitation of our approach is that it only considers heterogeneity with respect to a univariate quantity, i.e. the dimension of $V_{i}$ is $1$. If the treatment effect is potentially heterogeneous with respect to the full set of covariates $X_{i}$, i.e. the true model is $Y_{i}=\beta(X_{i})D_{i}+g(X_{i})+\epsilon_{i}$, equation (22) identifies $\mathbb{E}[\beta(X_{i})w(Z_{i},X_{i},v)|V_{i}=v]$ with the weight function
$w(Z_{i},X_{i},v)=\frac{(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}}{\mathbb{E}[(\mathbb{E}[D_{i}|Z_{i},X_{i}]-\mathbb{E}[D_{i}|X_{i}])^{2}|V_{i}=v]}.$
In particular, this is not just the projection $\mathbb{E}[\beta(X_{i})|V_{i}=v]$ of the true heterogeneous treatment effect on $V_{i}$.
A straightforward extension of our approach would be to use multivariate kernel smoothing to deal with multivariate $V_{i}$ and the results should carry over. However, note that the conditions on the convergence rates of the nuisance function estimators (Assumption 5) depend on the bandwidth $h$. In higher dimensions, they will depend on $h^{d}$, where $h$ is the dimension of $h$ and it will quickly become unrealistic to have machine learning methods that achieve these convergence rates.
We can also take one step back. In principle, we can replace the outer conditional expectations in (22) by arbitrary machine learning methods potentially also allowing for higher-dimensional $V_{i}$. A natural assumption might be that $V_{i}=X_{i}$ which is potentially high-dimensional, but that the treatment effect is only heterogeneous with respect to a small subset of the components of $X_{i}$, so one could use methods enforcing sparsity for the outer conditional expectation. If the machine learning method in the numerator allows for inference and the machine learning method in the denominator is consistent, one might still be able to perform inference for the heterogeneous treatment effect.
The aim of this work was to present a simple and ready-to-use method for inference in combination with user-chosen machine learning algorithms, but from a theoretical perspective, we did not strive for full generality. For the homogeneous case, estimators with better theoretical properties could probably be obtained using higher order influence functions [36] or triple sample splitting [32]. For the heterogeneous treatment effect – at least in the case of binary treatment and without endogeneity – minimax optimality was discussed by [29, 30]. Whereas this was not the focus of our work, it would be interesting to obtain similar results for the endogenous setting and allowing for continuous treatment as considered here.
@@ Acknowledgements
We are grateful to Malte Londschien for helpful discussions about weak instrumental variables. CS received funding from the Swiss National Science Foundation, grant no. 214865. ZG acknowledges financial support for visiting the Institute of Mathematical Research (FIM) at ETH Zurich. ZG was partly supported by NIH R01LM013614, NIH R01AG086379, and NSF DMS 2413107.
@@ Appendix A Technical Conditions for the Heterogeneous Treatment Effect
The following assumptions are needed to be able to estimate and perform inference for the heterogeneous treatment effect $\beta(v)$.
@@
    The density $p_{V}(\cdot)$ of $V_{i}$ with respect to Lebesgue measure satisfies that $p_{V}(v)\geq C$ for a constant $C>0$ independent of $N$.
@@
    $\sigma_{Q}^{2}(v)\asymp\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v]$.
@@
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=v\right]\lesssim\sigma_{Q}^{2}(v)$.
@@
    The function $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}(v)<C$.
@@
    The function $\rho(x)=\mathbb{E}[\epsilon_{i}^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(v))\mathbb{E}[\epsilon_{i}(D_{i}-\phi(X_{i}))(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(v))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
@@
    The function $\rho(x)=(\beta(x)-\beta(v))^{2}\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
@@
    The function $\rho(x)=\mathbb{E}[(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
@@
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}(f(Z_{i},X_{i})-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}/\sigma_{Q}^{2}(v)<C$.
@@
    The function $\rho(x)=\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|V_{i}=x]p_{V}(x)$ is twice differentiable for $x$ in the support of $V_{i}$ and there exists $C<\infty$ independent of $N$ such that $\|\rho^{\prime\prime}\|_{\infty}<C$.
@@
    $\frac{\mathbb{E}\left[\left|\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)\left(f(Z_{i},X_{i})-\phi(X_{i})\right)K\left(\frac{V_{i}-v}{h}\right)\right|^{2+\eta}\right]}{\sigma_{Q}(v)^{2+\eta}}\leq Ch.$
@@
    $\frac{\mathbb{E}\left[\left(\epsilon_{i}+(\beta(V_{i})-\beta(v))(D_{i}-\phi(X_{i}))\right)^{4}\left(f(Z_{i},X_{i})-\phi(X_{i})\right)^{4}K\left(\frac{V_{i}-v}{h}\right)^{4}\right]}{\sigma_{Q}^{4}(v)}\ll Nh^{2}.$
@@
    $\mathbb{E}\left[(D_{i}-\phi(X_{i}))^{4}K\left(\frac{V_{i}-v}{h}\right)^{2}\right]=O(h).$
@@
Note that all the quantities in Assumption 6 may depend on $N$ even if it is not reflected in the notation. Assertion 1 intuitively means that we need to have enough samples $i$ with $V_{i}$ close to $v$ to be able to estimate the heterogeneous treatment effect. Assertions 2 and 3 are needed in the case where the IV strength $\sigma_{Q}^{2}(v)$ is decreasing. Essentially they state that a decreasing $\sigma_{Q}^{2}(v)$ is attributed to $f(Z_{i},X_{i})$ being close to $\phi(X_{i})$. They are for example satisfied if $\mathbb{E}[\epsilon_{i}^{2}|Z_{i},X_{i}]$ and $\mathbb{E}[(D_{i}-\phi(X_{i}))^{2}|Z_{i},X_{i}]$ are both bounded away from zero and infinity independent of $N$. Assertions 4-11 are smoothness assumptions on $\beta(x)$, $p_{V}(x)$ and conditional moments of $\epsilon_{i}$, $D_{i}$, $f(Z_{i},X_{i})$ and $\phi(X_{i})$. These smoothness conditions need to hold independently of $N$. In view of Lemma 23 in Appendix D.5 (with the kernel $K$ potentially replaced by $K^{2+\eta}$ or $K^{4}$), assertions 12-14 are comparably weak and hold in particular, if the quantities inside the expectations satisfy similar smoothness conditions as assertions 4-11.
@@ C.1 Explicit Expression for the Robust Confidence Set
Using basic manipulations, we can derive explicit expressions for $\mathcal{C}_{N}^{\alpha}(v)$. For this, define the quantities
$A_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$
$B_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{D,i}^{k}R_{f,i}^{k}K\left(\frac{V_{i}-v}{h}\right)$
$C_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{Y,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$E_{N}$ $=\frac{2}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}R_{Y,i}^{k}R_{D,i}^{k}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}$
$F_{N}$ $=\frac{1}{Nh}\sum_{k=1}^{K}\sum_{i\in I_{k}}\left(R_{D,i}^{k}\right)^{2}\left(R_{f,i}^{k}\right)^{2}K\left(\frac{V_{i}-v}{h}\right)^{2}.$
@@
$\mathcal{C}_{N}^{\alpha}(v)=\left\{\gamma:(A_{N}-\gamma B_{N})^{2}\leq\frac{z_{1-\alpha/2}^{2}}{Nh}\left(C_{N}-\gamma E_{N}+\gamma^{2}F_{N}-h(A_{N}-\gamma B_{N})^{2}\right)\right\}$
or equivalently $\mathcal{C}_{N}^{\alpha}(v)=\left\{\gamma:R_{N}\gamma^{2}+S_{N}\gamma+T_{N}\leq 0\right\}$ with
@@
$\mathcal{C}_{N}^{\alpha}(v)=\begin{cases}\left[\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}>0\\ \mathbb{R}\setminus\left[\frac{-S_{N}+\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}},\frac{-S_{N}-\sqrt{S_{N}^{2}-4R_{N}T_{N}}}{2R_{N}}\right],&R_{N}<0.\end{cases}$ (27)
By construction, we have that $\hat{\beta}(v)=A_{N}/B_{N}\in\mathcal{C}_{N}^{\alpha}(v)$. Hence, in the case $R_{N}>0$, it will always hold that $S_{N}^{2}-4R_{N}T_{N}>0$. In the case $R_{N}<0$, it may happen that $S_{N}^{2}-4R_{N}T_{N}<0$, in which case, $\mathcal{C}_{N}^{\alpha}(v)=\mathbb{R}$.